A wire has a resistance of 280 Ω .

What will the wire's resistance be if it is stretched to be 4.5 times its original length but its volume stays the same?
Answer #1

Solution)

Given, R=280 ohms

Let length be L m

Cross-sectional area= A m^2

Now, after stretch length becomes 4.5 L

And , area becomes A1

Given,

Initial volume= Final Volume

L*A= 4.5L*A1

So, A1=A/4.5

We know,

R= rho*L/A =0.25 (at t=0)

And, R1=rho*(4.5 L/A/4.5)=9*rho*L/A =9*R =9*0.25= 2.25 ohms (Ans)

========